### v2

# Meta-Reasoning Improves Tool Use in Large Language Models

Abstract
External tools help large language models succeed at tasks where they would otherwise typically fail. In existing frameworks, choosing tools at test time relies on naive greedy decoding, regardless of whether the model has been fine-tuned on tool-annotated data or prompted with in-context examples. In contrast, we find that gathering and choosing among a suitable set of candidate tools has greater potential to lead to an optimal selection. We present Tool selection via meta-reasoning (Tecton), a two-phase system that first reasons over a task and outputs candidate tools using a custom fine-tuned language modelling head. Then, with the custom head disabled, it meta-reasons (i.e., it reasons over the previous reasoning process) to make a final choice. We show that Tecton results in substantial gains—both in-distribution and out-of-distribution—on a range of math reasoning datasets.
Meta-Reasoning Improves Tool Use in Large Language Models
Lisa Alazraki Imperial College London lisa.alazraki20@imperial.ac.uk                        Marek Rei Imperial College London marek.rei@imperial.ac.uk
1 Introduction
Figure 1: An overview of Tecton. In the reasoning phase, the system inspects the task and decodes a set of candidate tools, followed by argument insertion and evaluation of the tools via the Python interpreter. In the meta-reasoning phase, the model is asked to select the most useful tool, either by scoring multiple options (Tecton-score) or by continuing the generation given the decoded tools as hints (Tecton-generate).
Augmentation with external tools has proven effective at boosting the performance of large language models (LLMs) in knowledge-intensive tasks such as QA and math problem-solving Hao et al. (2023); Paranjape et al. (2023); Parisi et al. (2022); Schick et al. (2023). Tools are self-contained programs or APIs which the model can execute with chosen arguments as needed. To teach an LLM how to use tools, previous work adopts one of three main strategies: (1) tool demonstrations via in-context learning (ICL) Gao et al. (2023); Gupta and Kembhavi (2023); Hsieh et al. (2023); Surís et al. (2023), (2) full model fine-tuning on a dataset where text samples are interleaved with tool annotations Parisi et al. (2022); Schick et al. (2023); Tang et al. (2023); Patil et al. (2023), or (3) parameter-efficient fine-tuning (PEFT) on tool annotated data Hao et al. (2023); Qiao et al. (2024); Wang et al. (2024). Similarly to full fine-tuning, PEFT methods can teach LLMs a very large number of tools, while ICL is limited by the fixed size of the context window Hao et al. (2023); Patil et al. (2023). Although the fine-tuning paradigm in general binds the model to the set of tools learned during training, parameter-efficient learning reduces the cost of tuning, thus facilitating potential future extensions of the tool set. In contrast, adding further tools via a new round of full-model tuning incurs a significant computational cost Hao et al. (2023). We note that a further advantage of PEFT is that it tunes a handful of additional task-specific parameters, which can be selectively disabled to reinstate the frozen model and its original capabilities Ding et al. (2023); Han et al. (2024).
In existing literature, inference-time tool selection is made by greedily decoding the most likely tool Gao et al. (2023); Hao et al. (2023); Schick et al. (2023); Wang et al. (2024), regardless of whether the model has been fully fine-tuned, PEFT-tuned, or prompted in-context. In this work, we revisit the paradigm that selects tools solely based on their probability at decoding time. We propose an alternative, novel framework for Tool selection via meta-reasoning (Tecton) that gathers and chooses among a suitable set of candidate tools. We train a parameter-efficient language modelling head on tool-annotated data, similar to Hao et al. (2023), which can be switched on or off as needed. Tecton thus comprises two distinct phases: in the reasoning phase, it investigates a task and outputs candidate tools with the aid of the tuned LM head. Then, in the meta-reasoning phase, it uses the frozen LLM to re-examine the candidates and make a final decision. Fig. 1 illustrates the framework. We train and evaluate Tecton on math reasoning datasets, following established work on LLM tool calling Chen et al. (2024); Das et al. (2024); Gou et al. (2024). Among tasks that benefit from tools, math reasoning is particularly challenging, since it requires chains of multiple tool calls with errors that compound. This is evidenced by existing tool-augmented LLMs, which achieve the lowest performance on math reasoning when evaluated on multiple tasks Hao et al. (2023); Schick et al. (2023). In summary, our main contributions are:
  • We introduce Tecton, a novel two-phase framework that combines a custom fine-tuned head with a frozen LLM to improve tool use (Section 2).
  • We show that Tecton outperforms strong baselines in math reasoning tasks, both on in-distribution data and on unseen benchmarks (Section 3).
  • We enhance three popular math reasoning datasets to make them more challenging for current LLMs. We share our data and code at https://github.com/lisaalaz/tecton.
2 Method
2.1 Preliminaries
We augment the language modelling head of a base model with additional token embeddings $T$ to represent math operations, and train them via a standard language modelling objective. Once trained, $T$ comprises the tools available to the LLM for solving math problems. Prior work has shown the effectiveness of tuning additional tokens for both math tasks Hao et al. (2023); Wang et al. (2024) and general reasoning Goyal et al. (2024); Herel and Mikolov (2024).
Our preliminary experiments show that in cases where the LLM has failed to generate the correct tool by greedy sampling, this can usually be found among tokens that have only slightly lower probabilities (Figure 3). Over-sampling an appropriate set of candidate tools may thus be a better strategy than greedily decoding the most likely tool, provided this is combined with a reliable method for choosing among the candidates. To this end, we design a two-phase framework that leverages both the specialised, augmented LM head (reasoning phase) and the underlying generalist LLM (meta-reasoning phase). The rest of the section describes this in detail.
2.2 Reasoning Phase
[…]
2.3 Meta-Reasoning Phase
In this phase, we disable the custom-tuned head and let the underlying LLM analyse its previous reasoning process to choose among the candidate tools. Frozen LLMs have been used to self-evaluate and meta-reason over previous answers in existing literature Alazraki et al. (2023); Shinn et al. (2023); Yao et al. (2023a); Zeng et al. (2024b). We experiment with two ways of eliciting meta-reasoning: Tecton-score and Tecton-generate.
Tecton-score.
We join each candidate tool with the previous context, and present these as options for the LLM to score. We prefix each option with an uppercase letter label and select as the answer continuation the option whose label is assigned highest probability by the model, i.e., $\text{arg max}_{t_{i}\in\mathcal{V}_{\text{sub}}}\,p(t_{i}\mid t_{{}<i}\text{ {with} }t_{{}<i}\in\mathcal{V})$, where $\mathcal{V}_{\text{sub}}$ denotes a subset of the vocabulary containing only the uppercase letter tokens that are in the label set. In this setup, we limit the number of candidates to a maximum of four.
Tecton-generate.
We pass the candidate tools as hints and ask the model to generate an appropriate continuation of the answer. Here, the hints serve as mere guidance for the LLM (i.e., the model could choose to ignore all candidates and generate something different). In this version of the system, the model benefits from dynamically retrieved few-shot exemplars demonstrating the tools in the candidate set.
Figure 2: Averaged biased probability distributions over $n$ labels (for $n=\{1,2,3\}$), obtained on GSM8K-XL’s validation set. The dotted lines indicate the uniform averaged distribution that would be given by an unbiased model. Note that the label distribution is similarly skewed for FuncQA.
2.4 Bias Calibration
Upon running Tecton-score without recalibration of the label probabilities, we find that the validation results are poor. Visual inspection of the samples reveals that the model assigns highest likelihood to the same label in most instances, as shown in Fig. 2. This is consistent with Zheng et al. (2024)’s finding that LLMs are prone to selection bias in multiple-choice tasks. To solve a similar problem, Duarte et al. (2024) measure their model’s bias over the labels $A$, $B$, $C$, $D$ using a set of neutral samples where a uniform distribution would be expected, and subtract that bias from each label’s likelihood at inference time. Our math reasoning task does not lend itself to finding neutral samples, so we adopt a different strategy. Having created data samples of questions and options (by running the reasoning phase of Tecton on a validation set), we compute $n!$ permutations of the $n$ options while keeping the letter labels in the same position. We have the model score the labels for each permutation, and average over all permutations and all data samples to obtain an averaged biased distribution $B^{(n)}$. Since the number of options in our task is variable, we run this process independently for data samples with $n=2$, $n=3$ and $n=4$ labels. At inference time, given a set of labels $L^{(n)}$ of size $n$, we retrieve the corresponding $B^{(n)}$ and compute the calibrated probability $\hat{p}_{i}$ of each label $l_{i}$ in the set as
$\hat{p}_{i}=p_{i}+\frac{1}{n}-{B^{(n)}}_{i}$
where $p_{i}$ is the probability assigned by the model to label $l_{i}$ for the current sample, ${B^{(n)}}_{i}$ is the pre-computed biased probability of label $l_{i}$, and $n$ is the total number of labels.
2.5 Retrieval of Tool Demonstrations
To aid answer generation in Tecton-generate, we dynamically retrieve and add to the context few-shot exemplars demonstrating the candidate tools. We create the retrieval pool by extracting training samples and collecting candidate tools for each, by running the reasoning phase of Tecton. We construct each exemplar to simulate the inference task, as follows: (1) we append to the sample its set of candidate tools, and (2) we append to it the golden answer demonstrating how the correct tool is used to obtain the final solution. At inference time, we retrieve only the exemplars whose golden answers contain the tools currently in the candidate set. It is worth noting that dynamic retrieval was not included in Tecton-score as it did not improve validation performance.
3 Experiments
3.1 Experimental Setup
Our system is model-agnostic and can be applied to any open-weights LLM. Here, we use Llama 3 8B Instruct Dubey et al. (2024) (henceforth referred to as Llama 3) as the base model in all experiments. Implementation details and hyperparameters are given in Appendix B.
3.2 Datasets
We train and evaluate Tecton on GSM8K-XL Cobbe et al. (2021); Hao et al. (2023) and FuncQA Hao et al. (2023). The test set of the latter is comprised of two distinct subsets: a ‘one-hop’ corpus containing problems solvable with one single operation (FuncQA-OH), and a ‘multi-hop’ one requiring multiple operations (FuncQA-MH). For GSM8K-XL we tune four additional tokens corresponding to the four basic operations, and extend these to 13 in the case of FuncQA. The complete set of tools for each dataset is shown in Appendix B. Additionally, we evaluate on out-of-distribution datasets that were not observed during fine-tuning. For this purpose we choose a range of math reasoning datasets: ASDiv Miao et al. (2020), MAWPS Koncel-Kedziorski et al. (2016) and SVAMP Patel et al. (2021). These were found by Ott et al. (2023) to be OOD with respect to GSM8K: not only do they display minimal n-gram overlap, but they also require reasoning chains of different average lengths from GSM8K to be solved. These datasets have been used to evaluate LLMs in previous works, both with and without the aid of tools Kojima et al. (2022); Schick et al. (2023). Following a strategy similar to the one used by Hao et al. (2023) for constructing GSM8K-XL, we magnify the numbers in these datasets to make them challenging for current LLMs (the enhancement process is described in Appendix D). We thus obtain ASDiv-XL, MAWPS-XL, and SVAMP-XL. We use these datasets to test models tuned on GSM8K-XL.
3.3 Baselines
We implement recent tool-augmented models as baselines: Trice Qiao et al. (2024) and ToolkenGPT Hao et al. (2023). These share a parameter-efficient approach with Tecton (see Appendix C for details). Despite their relatively low computational cost, they have been shown to outperform strong systems: Trice paired with Alpaca Taori et al. (2023), ChatGLM Zeng et al. (2024a) and Vicuna Zheng et al. (2023) surpasses the much larger GPT-3.5 as well as tool learning via supervised fine-tuning. In addition to outperforming GPT-3.5, ToolkenGPT paired with LLaMA Touvron et al. (2023) is more accurate than ReAct Yao et al. (2023b). It should also be noted that LLMs are increasingly able to solve math problems and perform difficult arithmetic without tools, as shown by the high accuracy ($79.6\%$) achieved by Llama 3 on the non-enhanced version of GSM8K Dubey et al. (2024). Hence we additionally compare against a vanilla version of Llama 3 as well as Chain-of-Thought (CoT) prompting Wei et al. (2022) with exemplars extracted from the train set.
In-distribution Out-of-distribution
FuncQA-OH FuncQA-MH GSM8K-XL ASDiv-XL MAWPS-XL SVAMP-XL
  Llama 3 10.0 2.9 13.0 25.8 26.2 27.8
$+$ CoT 25.0 5.9 37.3 40.6 58.7 51.9
  $+$ Trice - - 43.5 52.2 71.2 49.6
$+$ ToolkenGPT 65.0 10.3 48.4 45.3 68.3 60.4
  $+$ Tecton-score 66.7 17.6 50.7 53.6 76.9 62.2
$+$ Tecton-generate 70.0 20.6 45.8 55.3 77.4 66.7
Table 1: Accuracies on math reasoning datasets measured via exact match of the result rounded to 2 decimal places. We do not tune Trice on FuncQA as this dataset lacks the necessary annotations for RLEF training. Note that the separation between in-distribution and out-of-distribution data does not apply to the vanilla version of Llama 3.
In-distribution Out-of-distribution
FuncQA-OH FuncQA-MH GSM8K-XL ASDiv-XL MAWPS-XL SVAMP-XL
  Tecton-score – bias calibration 58.3 8.8 49.3 52.2 75.0 61.9
  Tecton-generate – dynamic exemplar retrieval 58.3 13.2 43.5 50.8 71.2 62.9
Table 2: Results of ablating components of Tecton-score and Tecton-generate, on in-distribution and out-of-distribution data.
3.4 Results
Table 1 shows Tecton’s gains on math reasoning. Both versions of the system achieve scores above baselines for in-distribution and out-of-distribution-data, with the exception of Tecton-generate on GSM8K-XL, whose performance is slightly below that of ToolkenGPT.
In-distribution performance.
On GSM8K-XL, our best implementation scores $7.2$ percentage points above Trice and $2.3$ above ToolkenGPT. When evaluated on in-distribution data, Tecton’s most significant gains are on FuncQA: the generate setting doubles the performance of ToolkenGPT in the challenging multi-hop task, reaching $20.6\%$ accuracy. For context, Llama 3 can only solve $2.9\%$ of the dataset and CoT only raises performance to $5.9\%$.
Out-of-distribution performance.
On OOD datasets, Tecton’s advantage is substantial: on average, Tecton-generate gains $8.5$ percentage points over ToolkenGPT and $8.8$ over Trice. Tecton-score achieves $6.2$ and $6.6$ above the two baselines, respectively. This highlights the adaptability of the method to unseen data.
3.5 Ablations
To gain more insight into these results, we perform an ablation study on the meta-reasoning phase of Tecton, shown in Table 2. We ablate bias calibration from Tecton-score and dynamic exemplar retrieval from Tecton-generate. On average, Tecton-score’s ablated accuracy is $6.2$ percentage points lower than the non-ablated version on in-distribution data, and $1.2$ on unseen datasets. Additionally, Tecton-generate’s average performance drops by $7.2$ on in-distribution data and $4.9$ on OOD data. The most significant performance loss is on FuncQA: the ablated versions of Tecton-score and Tecton-generate see a decrease of $8.4$ and $11.7$ percentage points, respectively, on the one-hop test set. They also drop by $8.8$ and $7.4$, respectively, on the multi-hop split. On GSM8K, the decline is less severe: $-1.4$ for Tecton-score and $-2.3$ for Tecton-generate. Tecton-score’s decrease after ablation is similarly modest on OOD datasets, while Tecton-generate’s is more significant: its ablated performance drops by $6.2$ points on MAWPS, $4.5$ on ASDiv and $3.8$ on SVAMP. This highlights the benefit of dynamic exemplar retrieval during the meta-reasoning phase of Tecton-generate. Despite the performance decline, we find that both ablated versions of Tecton outperform CoT on all datasets. They also consistently surpass ToolkenGPT’s accuracy on OOD datasets, and either match or outperform Trice, with the sole exception of Tecton-generate on ASDiv-XL.
4 Conclusion
We introduce Tecton, a novel two-phase framework that first samples a set of candidate tools and then selects the optimal candidate via meta-reasoning. We implement two versions of the system and find that both achieve superior performance on math reasoning datasets, surpassing our strongest baseline by $\sim 4\%$ on in-distribution data and $\sim 9\%$ on unseen benchmarks, on average. These results confirm our hypothesis that a specialized, custom-tuned framework and a generalist pre-trained model can work together to improve tool use in challenging tasks.
Limitations
This paper solely focuses on math reasoning tasks. While this is consistent with established literature, there are other domains (e.g., knowledge-intensive QA, virtual environment navigation) that can benefit from the use of tools. Future work can investigate a wider range of tasks.
Ethical Considerations
We have verified that all datasets and software utilized in this paper allow for their use, distribution and modification. Our non-commercial purpose is consistent with all licenses. The distribution of our code and data is accompanied by the licenses and credits to the original authors.
Acknowledgments
The authors would like to thank Joe Stacey for his insightful comments on the first draft of this paper.
Appendix A Preliminary Experiments
A.1 Top-k Decoding
We run ToolkenGPT (in its original implementation based on the first version of LLaMA) on the validation set of FuncQA. We find that in samples where the system has generated an incorrect tool by greedy sampling, the correct one—complete with correct arguments—is among the top five most likely tokens in over $60\%$ of cases. These include samples where the correct tool can be found by searching the top tokens at a different position from the one that has produced the incorrect tool. Overall, the system decodes the correct tool (regardless of the arguments it generates for it) in 76.9% of samples; this rises by over $10\%$ to $87.2\%$ when we consider the top $k=5$ tokens at each decoding step (Fig. 3). Further increasing $k$ to $10$ incurs a higher computational cost without raising the proportion of decoded golden tools.
A.2 Temperature Sampling
We experiment with temperature sampling and find that it is not an optimal strategy to gather candidate tools, as lower temperatures do not lead to enough diversity while higher ones generate irrelevant tools.
A.3 Tool Selection by Self-Consistency
We select tools by self-consistency on FuncQA’s validation set. We try choosing the tool that is most represented in the candidate set both before and after argument insertion. In both cases, We find that tool choice by self-consistency performs poorly (over 6% below ToolkenGPT in the best case).
Figure 3: Proportion of golden tools across the FuncQA validation set, for k=1, k=5 and k=10, where k is the number of top-probability tokens at each decoding step.
Appendix B Model Implementation and Training
We do all training and inference on a single NVIDIA Tesla V100 GPU.
B.1 Tools
GSM8K-XL and FuncQA Hao et al. (2023) are annotated with four and 13 tools respectively, each representing a math operation. We illustrate these in Table 3. Tools are trained as additional tokens added to the standard language modelling head of Llama 3, which comprises 128,256 token representations. Therefore, we extend these representations to 128,269 in the case of FuncQA and 128,260 for GSM8K-XL.
B.2 Details of the Training Process
The LM head of Tecton consists of the standard head of Llama 3 8B Instruct Dubey et al. (2024) concatenated with an additional linear layer of size $embedding\_dimension\times number\_of\_tools$. The tool token embeddings are randomly initialized and trained on math reasoning QA pairs from GSM8K-XL and FuncQA, following a standard language modelling objective. Both datasets are annotated with the token positions at which each tool should be generated. Note that the original datasets made available by Hao et al. (2023) are annotated according to SentencePiece tokenization Kudo and Richardson (2018). We edit the annotations to be compatible with Llama 3’s Byte-Pair Encoding tokenizer Sennrich et al. (2016). We tune two distinct sets of special tokens, one on each dataset, as each requires a different set of tools as shown in Table 3. We train at half precision (FP16) for ten epochs, using learning rates {$1$e44-4- 4, $1$e33-3- 3}, batch size 1, and saving checkpoints at each epoch. We select the best checkpoint by measuring performance on a validation set.
Table 4 gives an overview of our training, validation, and testing data. It should be noted that GSM8K-XL’s training and validation sets coincide with those of GSM8K, as only the test portion of the dataset was enhanced by Hao et al. (2023). Table 5 reports the hyperparameter combinations of our chosen checkpoints.
Task Tools
GSM8K-XL <add> <subtract> <multiply> <divide>
FuncQA <add> <subtract> <multiply> <divide> <power> <sqrt> <log> <ln> <lcm> <gcd> <remainder> <choose> <permutate>
Table 3: Math reasoning datasets and their corresponding tools.
Dataset Train Validation Test
GSM8K-XL 5054 1000 568
FuncQA 611 39 128
ASDiv-XL N/A N/A 360
MAWPS-XL N/A N/A 416
SVAMP-XL N/A N/A 270
Table 4: Sizes of our training and testing datasets. Note that we use ASDiv-XL, MAWPS-XL and SVAMP-XL only for testing. For FuncQA the reported test data size includes both splits—FuncQA-OH and FuncQA-MH.
Dataset Epoch Learning Rate
GSM8K-XL 9 1e33-3- 3
FuncQA 4 1e44-4- 4
Table 5: Hyperparameter combinations of the chosen model checkpoints for GSM8K-XL and FuncQA.
B.3 Inference-time Hyperparameters
At inference, we decode with temperature $t=0$, $p=0.95$, $k=5$. In the reasoning phase, we apply logit bias to the tool tokens to promote their generation. We use logit bias $3.0$ for GSM8K-XL and $4.0$ for FuncQA. Since our generation strategy is deterministic, all our accuracies are reported on a single run, rounded to 1 decimal place.
Appendix C Baselines Implementation
All baselines are built upon Llama 3 8B Instruct.
Llama 3 8B Instruct.
We measure our base model’s performance in the zero-shot setting. We experiment with CoT zero-shot prompting (prepending to the question the instruction Let’s think step by step) but find that just using the raw question as input results in better performance.
Chain-of-Thought (CoT).
We extract from the training set six pairs of questions and answers demonstrating maths operations and their use. For comparability, we use the same exemplars as in the implementation of ToolkenGPT and Tecton. Note that Llama 3 8B Instruct achieves $79.6\%$ accuracy Dubey et al. (2024) on the non-enhanced version of GSM8K when prompted in few-shot CoT fashion. Our experiments show that it performs significantly worse ($37.3\%$) on GSM8K-XL, highlighting the difficulty of the enhanced dataset.
Trice.
Trice is a two-stage system where instruction tuning is followed by reinforcement learning from environmental feedback (RLEF), leveraging LoRA at both stages. We train Trice using the same hyperparameters as in the original implementation, with the exception of the batch size in the first phase of training, reduced to 128 with 8 gradient accumulation steps due to the memory limitations of our hardware. Since we train on a single dataset (GSM8K-XL), our implementation corresponds to the setup referred to as Trice-split in the original paper. Note that Qiao et al. (2024) implement Trice on top of Alpaca Taori et al. (2023), ChatGLM Zeng et al. (2024a) and Vicuna Zheng et al. (2023). For comparability with Tecton and the other baselines, here we use Llama 3 as the base model. We thus adjust Trice’s prompt templates and special tokens to be consistent with Llama 3’s model card [^1] https://www.llama.com/docs/model-cards-and-prompt-formats/meta-llama-3.
ToolkenGPT.
Like Tecton, ToolkenGPT augments the output matrix of an LLM with additional special tokens, each representing a tool. Only the additional tokens are tuned while the rest of the weights remain frozen.We implement ToolkenGPT with Llama 3 and train it on the same annotated datasets as we train Tecton, using the same sets of tools and the same hyperparameter combinations for direct comparability.
Test set Original XL version
ASDiv 618 360
MAWPS 505 416
SVAMP 299 270
Table 6: Test set sizes of our OOD math reasoning datasets, in their original version and in the XL version enhanced via an automated process.
Appendix D Out-of-distribution Datasets
We enhance the test sets of ASDiv, MAWPS and SVAMP and obtain ‘XL’ versions of each. Our goal is to replace the numbers in each test sample with larger-magnitude ones in an automated manner, yet ensuring that the new numbers in each question correctly map to a new numerical answer. To this end, we use datasets that have been annotated with the golden sequence of operations. Since the original versions of ASDiv and SVAMP do not contain such annotations, we use the versions that are part of the Lila benchmark Mishra et al. (2022).
Firstly, we extract all the numbers from the chain of operations using regular expressions. Secondly, we search for these numbers in the question text (we use the num2words library [^2] https://github.com/savoirfairelinux/num2words to include numbers expressed as words) and replace them with random numbers in the interval $[-10^{5},10^{5}]$. We replace the numbers in both the text and the annotated chain of operations. Note that we replace negative numbers with negative numbers and positive numbers with positive numbers. We also take care to substitute integers with integers and floats with floats. If we cannot match any of the numbers in the operations to the text, we discard that sample. Lastly, we evaluate the new chain of operations with the substituted numbers and store the result as the new golden answer for that data point.
This process results in datasets containing questions with the same structure and wording as the original ones, but with the numbers greatly magnified. This provides a real challenge for contemporary large language models, which can otherwise easily solve small-number arithmetic.
As our automated process discards any data samples that it is unable to process, our final test sets are typically smaller than their original counterparts. Table 6 reports the sizes of the original and modified test sets.
Appendix E Prompts
[!t] Prompt 1: Tecton-score meta-reasoning prompt In this task, you must select the correct and most plausible continuation for the answer in the text below. You should use your common sense and logical and mathematical abilities. You should directly answer by choosing the correct option. Output only the letter corresponding to the correct option.
[FIXED EXEMPLARS]
Text to continue:
Question: [QUESTION]
Answer: [PARTIAL ANSWER]
Possible answer continuations:
[OPTIONS]
The correct continuation is:
[!t] Prompt 2: Tecton-generate meta-reasoning prompt Complete the answers below. In square brackets you will find some hints showing the possible math operations. You may use one of these hints if you think it is correct.
[DYNAMIC EXEMPLARS]
Question: [QUESTION]
Answer: [HINTS] [PARTIAL ANSWER]
In the reasoning phase, we prompt Tecton to generate an answer to a math reasoning question step by step, aided by in-context exemplars. Prompts E and E illustrate how we elicit tool choice in the meta-reasoning phase, for Tecton-score and Tecton-generate respectively. In both cases, we show the model in-context exemplars, followed by the current question and the current partial answer. The latter consists of the lines of text generated during the previous iterations of Tecton, if any. In Prompt E, the placeholder [FIXED EXEMPLARS] is replaced with six in-context exemplars extracted from the training set. The number of options in each exemplar matches the number of available options in the current sample. In Prompt E, we have [DYNAMIC EXEMPLARS] that are retrieved to demonstrate the tools currently in the candidate set. Here, the partial answer also includes the tokens generated during the current iteration, up to the sequence position where the first tool is found among the $k$ top probability tokens. The hints are prepended to the partial answer in this setting.
Appendix F Efficiency
We compare the efficiency of our method with the fine-tuned baselines—Trice and ToolkenGPT.
F.1 Comparison with Trice
Tecton is more efficient than Trice, as it only requires tuning additional embeddings in the output layer, with minimal backpropagation. In contrast, Trice performs two rounds of finetuning each time updating LoRA modules throughout the model.
F.2 Comparison with ToolkenGPT
We train Tecton in the same way as the ToolkenGPT baseline, by only updating the additional token embeddings in the output layer. At inference, Tecton requires generating additional text for each line in the problem. Note that for Tecton-score this overhead is negligible, as we generate only one additional token per line.